OpenStudy (justme):

please help 2(2x+1)=8

OpenStudy (anonymous):

So 4x + 2 = 8 so 4x = 8-2 = 6 x = 6/4 = 1.5

OpenStudy (anonymous):

it also depends on what u want done what u learnin

OpenStudy (justme):

quadriatic functions

OpenStudy (anonymous):

ok so X=-b+-square root b^2-4ac over 2a

OpenStudy (justme):

can u factor 4x-6=0 instead of using the quad. formula?

OpenStudy (anonymous):

hold on let me check

OpenStudy (anonymous):

i dont think u can the book dosent say anything about that

OpenStudy (justme):

ok thanks!

OpenStudy (anonymous):

wait there are 5 ways to solve the quadratic formula wold u like to hear them

OpenStudy (justme):

yes sure

OpenStudy (anonymous):

first theres the graphing way only if u have a graping calculater handy

OpenStudy (justme):

theres no graphing in this section of our book

OpenStudy (anonymous):

two thers square roots with no x term three factoring four compleating the square use if the x^2 term is one ,but you cant factor the thing esey lastly the quadratic formula

OpenStudy (anonymous):

sorry i couldnt be of more help

OpenStudy (justme):

no problem. its a take home test. i can figure this will be hard. thanks anyway.

OpenStudy (anonymous):

hahaha sorry

OpenStudy (radar):

Justme, the equation you submitted was not a quadratic, as it did not have a variable to a power such as:\[x ^{2}\] this is a linear equation.2(2x+1)=8 Remove the parenthesis by distributing the 2 getting 4x+2=8 then subtract 2 from both sides getting: 4x=8-2=6 We now have the value for 4x as 4x=6. However, you want the value of x. Divide both sides by 4 getting: (4/4)x=6/4 1x=3/2 or x=3/2=1 1/2